OpenStudy (saifoo.khan):

Jon's new mini computer monitor has a 20 cm by 24 cm LCD screen with a plastic trim border. The total area of the monitor and border is 672 cm2. How wide is the trim around the monitor?

OpenStudy (anonymous):

-36 = -1 + 5x

OpenStudy (saifoo.khan):

How?

OpenStudy (anonymous):

that seems extremyl wrong to me

OpenStudy (anonymous):

this is basic math..... you can do it saif

OpenStudy (saifoo.khan):

im getting 4. how about it?

OpenStudy (anonymous):

of course your getting 4

OpenStudy (saifoo.khan):

is it correct?

OpenStudy (saifoo.khan):

(24+w)(20+w)=672 480+44w+w^2=672 w^2 + 44w - 192 = 0 (w+48)(w-4) = 0 w = -48 or w = 4 we will neglect the negative value, So, w = 4